# Basic representations of genus zero nonabelian Hodge spaces

### v2

@@ 2.3. Diagrams
    If $a_{I}=a_{J}=\infty$ then $B_{IJ}=B^{\infty}_{IJ}$,
@@
    If $a_{I}=a_{J}\neq\infty$ then $B_{IJ}=B_{JI}=B^{\infty}_{IJ}-\alpha_{I}\beta_{J}-\alpha_{J}\beta_{I}$,
@@
    If $a_{I}=\infty$ and $a_{J}\neq\infty$ then $B_{IJ}=\beta_{I}(\alpha_{J}+\beta_{J})$,
@@
    If $a_{I}\neq\infty$ and $a_{J}=\infty$ then $B_{IJ}=B_{JI}=\beta_{J}(\alpha_{I}+\beta_{I})$,
  5. (5)
    If $a_{I}\neq\infty,\;a_{J}\neq\infty$ and $a_{I}\neq a_{J}$ then $B_{IJ}=B_{JI}=0$.
@@
The core diagram $\Gamma_{c}(\bm{\breve{\Theta}})$ is entirely determined by the fission data of $\bm{\breve{\Theta}}$ together with the slopes of all Stokes circles at finite distance, because these data determine the number $A_{IJ}=\operatorname{Irr}(\operatorname{Hom}(I,J))$. In general, the diagram contains strictly less information, because $A_{IJ}$ is a sum of different contributions, coming from the levels of the common parts of $I,J$ and the fission exponent $f_{I,J}$. As we shall review below, the interest of the (core) diagram is that, unlike the fission data, it is invariant under symplectic transformations
2.4. Fourier transform and symplectic transformations
We now describe how the $\operatorname{SL}_{2}(\mathbb{C})$ action on connections acts on their formal data and fission data.
2.4.1. Fourier transform
The Fourier transform is the automorphism of the Weyl algebra $A_{1}=\mathbb{C}[z]\langle\partial_{z}\rangle$ defined by $z\mapsto-\partial_{z},\;\partial_{z}\mapsto z$. It induces a transformation on modules on the Weyl algebra: the Fourier-Laplace transform $F\cdot M$ of a $\mathbb{C}[z]\langle\partial_{z}\rangle$-module $M$ is the $\mathbb{C}[z]\langle\partial_{z}\rangle$-module obtained from $M$ by setting: $z\mapsto-\partial_{z},\;\partial_{z}\mapsto z$. The Fourier transform induces a transformation on the set of isomorphism classes of irreducible connections (if we exclude the case of the trivial rank one connection on $\mathbb{P}^{1}$).
@@ 2.4.1. Fourier transform
There exists a bijection, that we also denote by $F$, from the set of all modified formal data to itself, such that if $(E,\nabla)$ is a irreducible algebraic connection on a Zariski open subset of $\mathbb{P}^{1}$, $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ its modified formal data, $(E^{\prime},\nabla^{\prime})$ is the Fourier transform of $(E,\nabla)$ and $(\bm{\breve{\Theta}^{\prime}},\bm{\breve{\mathcal{C}}}^{\prime})$ its modified formal data, the following diagram commutes:
${(E,\nabla)}$${(E^{\prime},\nabla^{\prime})}$${(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})}$${(\bm{\breve{\Theta}^{\prime}},\bm{\breve{\mathcal{C}}}^{\prime})}$$\scriptstyle{F}$$\scriptstyle{F}$
This map $F$ is the formal Fourier transform. It acts independently on the Stokes circles: if we use the notation $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})=\sum_{I\in N}(I,\breve{\mathcal{C}}_{I})$ where $N$ is the set of active circles of $\bm{\breve{\Theta}}$ we have
$F\cdot\sum_{I\in N}(I,\breve{\mathcal{C}}_{I})=\sum_{I\in N}F\cdot(I,\breve{\mathcal{C}}_{I}).$
@@
This bijection is actually given by a Legendre transform. It can be computed explicitly, and takes slightly different forms depending on whether the Stokes circles are at finite distance or at infinity, and in the latter case on whether its slope is greater of less than 1. The next proposition describes the ramification orders, slopes and levels of the formal Fourier transform $F\cdot I$ of any Stokes circle $I$ as a function of those of $I$, as well as the transformation of the conjugacy classes of formal monodromies.
@@
Let $I$ be a Stokes circle, and $\breve{\mathcal{C}}_{I}\in\operatorname{GL}_{n}(\mathbb{C})$ a conjugacy class. Let $\breve{\mathcal{C}}_{F\cdot I}\in\operatorname{GL}_{n}(\mathbb{C})$ be the conjugacy class such that $F\cdot(I,\breve{\mathcal{C}}_{I})=(F\cdot I,\breve{\mathcal{C}}_{F\cdot I})$. We have the following:
@@
    If $I$ i.e. of the form $\langle\alpha z\rangle_{\infty}$, with $\alpha\in\mathbb{C}$, then $F\cdot I=\langle 0\rangle_{\alpha}$, and $\breve{\mathcal{C}}_{F\cdot I}=-\breve{\mathcal{C}}_{I}$.
@@
    If $I$ is of slope $\leq 1$ at infinity, of the form $\langle\alpha z+q\rangle_{\infty}$, with $\alpha\in\mathbb{C}$, and $q\neq 0$ of slope $<1$ then $F\cdot I$ is of the form $F\cdot I=\langle\widetilde{q}\rangle_{\alpha}$, with $\operatorname{Irr}(\tilde{q})=s,\operatorname{ram}(\tilde{q})=r-s$, $\operatorname{Levels}(\tilde{q})=\frac{r}{r-s}\operatorname{Levels}(q)$. Also $\breve{\mathcal{C}}_{F\cdot I}=(-1)^{s}\breve{\mathcal{C}}_{I}$.
@@
    If $I$ is of slope >1 at infinity, with $\operatorname{ram}I=r$, $\operatorname{Irr}(I)=s$, then $F\cdot I$ satisifies $\operatorname{ram}(F\cdot I)=s-r$, $\operatorname{Irr}(F\cdot I)=s$, and $\operatorname{Levels}(F\cdot I)=\frac{r}{s-r}\operatorname{Levels}(I)$. Also $\breve{\mathcal{C}}_{F\cdot I}=(-1)^{s}\breve{\mathcal{C}}_{I}$.
@@
    If $I=\langle q\rangle_{a}$ is an irregular circle at finite distance for $a\in\mathbb{C}=\mathbb{P}^{1}\smallsetminus\{\infty\}$, with $q\neq 0$, $\operatorname{ram}(q)=r$, and $\operatorname{Irr}(q)=s$, then $F\cdot I$ is a circle of slope $\leq 1$, at infinity, of the form $\langle-az+\tilde{q}\rangle_{\infty}$, with $\operatorname{ram}(\tilde{q})=r+s,\operatorname{Irr}(\tilde{q})=s$, and $\operatorname{Levels}(\tilde{q})=\frac{r}{r+s}\operatorname{Levels}(q)$. Also $\breve{\mathcal{C}}_{F\cdot I}=(-1)^{s}\breve{\mathcal{C}}_{I}$.
@@
    If $I=\langle 0\rangle_{a}$ is a tame circle at finite distance, then $F\cdot I=\langle-az\rangle_{\infty}$, and $\breve{\mathcal{C}}_{F\cdot I}=-\breve{\mathcal{C}}_{I}$.
@@
All statements except those describing the sets of levels are already contained in the works [42, 37] establishing the stationary phase formula. For the levels, some of the statements are proven in [30] using results on singularities of curves. Let us discuss how to derive them from [18]. In the case where $I$ is at infinity and of slope $>1$, the statement for the levels can be deduced quite directly from Proposition 4.8 and the proof of Theorem 4.9 of [18]. Indeed, the proposition implies that implies that all possible exponents of $F\cdot I$ belong to a set determined by the levels of $I$, and that if $k$ is a level of $I$ then $kr/(r+s)$ is an exponent of $F\cdot I$ (because the coefficient of this exponent is nonzero), and the proof of Theorem 4.9 there establishes that these properties satisfied by the possible exponents of $F\cdot I$ imply that the exponents of $F\cdot I$ of the form $kr/(r+s)$ with $k\in\operatorname{Levels}(I)$ are the only ones where the sequence of common denominators of exponents of $F\cdot I$ strictly decreases, which precisely means that they are the levels of $F\cdot I$. The other cases will be completely similar, the only difference being the slope of $F\cdot I$. ∎
To fully determine the fission data of the Fourier transform, we must also describe the fission exponents between distinct Stokes circles of the Fourier transform which are at the same singularities. This can be explicitly done:
@@
Let $I,J$ be two distinct circles. There are the following possibilities:
@@
    • If $I$ and $J$ have a common part, denoting $k$ its slope, we have $f_{F\cdot I,F\cdot J}=\frac{1}{k+1}f_{I,J}$
@@
    If $\pi(I)=\pi(J)=\infty$, $I,J$ are both of slope $\leq 1$ and are of the form $I=\langle\alpha z+q\rangle$, $J=\langle\alpha z+q^{\prime}\rangle$ with $\alpha\in\mathbb{C}$, and $q,q^{\prime}$ of slope $<1$, then $\pi(F\cdot I)=\pi(F\cdot J)=\alpha$ and
    • If $q$ and $q^{\prime}$ have no common part and respective slopes $k,l$, then $f_{F\cdot I,F\cdot J}=\max(\frac{k}{1-k},\frac{l}{1-l})$.
    • If $q$ and $q^{\prime}$ have a common part $q_{c}$, denoting $k$ its slope, we have $f_{F\cdot I,F\cdot J}=\frac{k}{1-k}f_{q,q^{\prime}}$.
@@
    If $\pi(I)=\pi(J)=\infty$, and $I,J$ are both of slope $>1$, then $\pi(F\cdot I)=\pi(F\cdot J)=\infty$ and
    • If $I$ and $J$ have no common part, then denoting $k=\operatorname{slope}(I)$, $l=\operatorname{slope}(J)$, we have $f_{I,J}=\max(\frac{k}{k-1},\frac{l}{l-1})$.
    • If $I$ and $J$ have a common part, let $k$ denote its slope, then $f_{F\cdot I,F\cdot J}=\frac{k}{k-1}f_{I,J}$
@@
All statements where there is no common part involved follow directly from the previous proposition. For the statements in the cases where there is a common part, in the case where $I,J$ are at infinity and of slope $>1$ this is established in the proof of Theorem 5.9 of [18] (the word ‘fission exponent’ wasn’t used there: in terms of the terminology used there it corresponds to the highest number among the exponents of the leading term of the different parts of $F\cdot I$ and $F\cdot J$). The other cases will be similar. ∎
2.4.2. General symplectic $\operatorname{SL}_{2}(\mathbb{C})$ transformations
The Fourier-Laplace transform is part of a larger group of transformations acting on modules over the Weyl algebra. Indeed, to any matrix
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
in $\operatorname{SL}_{2}(\mathbb{C})$, we can associate an automorphism of the Weyl algebra $A_{1}=\mathbb{C}[z]\langle\partial_{z}\rangle$ given by $z\mapsto az+b\partial_{z},\partial_{z}\mapsto cz+d\partial_{z}$. This induces an action of the group $\operatorname{SL}_{2}(\mathbb{C})$ of symplectic transformations on modules over the Weyl algebra (see [37]).
@@ 2.4.2. General symplectic $\operatorname{SL}_{2}(\mathbb{C})$ transformations
    The Fourier-Laplace transform $F$, corresponding to the matrix $\begin{pmatrix}0&1\\ -1&0\end{pmatrix};$
@@
    Twists at infinity $T_{\lambda}$, for $\lambda\in\mathbb{C}$, corresponding to the matrix $\begin{pmatrix}1&\lambda\\ 0&1\end{pmatrix};$
@@
    Scalings $S_{\lambda}$, for $\lambda\in\mathbb{C}^{*}$, corresponding to the matrix $\begin{pmatrix}\lambda^{-1}&0\\ 0&\lambda\end{pmatrix}.$
The geometric interpretation of twists and scalings is the following. The twist $T_{\lambda}$ corresponds to taking the tensor product with the rank one module $(\mathbb{C}[z],\partial_{z}+\lambda z)$, and the scaling $S_{\lambda}$ corresponds to do the change of variable $z\mapsto z/\lambda$ on $\mathbb{P}^{1}$.
As for the Fourier transform, any element of $\operatorname{SL}_{2}(\mathbb{C})$ induces a transformation on modules over the Weyl algebra, and on irreducible connections on Zariski open subsets of the affine line. Moreover, there exists again a corresponding formal transformation (that we will also denote by $A$) on modified formal data with a similar structure. The action of twists and scalings on formal data and fission data is much simpler than for the Fourier transform:
@@
Let $I$ be a Stokes circle, and $\breve{\mathcal{C}}_{I}\in\operatorname{GL}_{n}(\mathbb{C})$ a conjugacy class. Let $\breve{\mathcal{C}}_{T_{\lambda}\cdot I}\in\operatorname{GL}_{n}(\mathbb{C})$ be the conjugacy class such that $T_{\lambda}\cdot(I,\breve{\mathcal{C}}_{I})=(T_{\lambda}\cdot I,\breve{\mathcal{C}}_{T_{\lambda}\cdot I})$. We have the following:
@@
    If $\pi(I)=\infty$, then, $\pi(T_{\lambda}\cdot I)=\infty$, and if $I=\langle q\rangle_{\infty}$ we have $T_{\lambda}\cdot I=\langle q-\frac{\lambda}{2}z^{2}\rangle_{\infty}$. In particular $T_{\lambda}\cdot I$ has the same ramification order and set of levels as $I$, $\operatorname{slope}(T_{\lambda}\cdot I)=\max(\operatorname{slope}(\langle q\rangle),\operatorname{slope}(\langle q-\frac{\lambda}{2}z^{2}\rangle)$, for $\lambda\in\mathbb{C}$.
@@
    Otherwise if $\pi(I)\neq\infty$, $T_{\lambda}\cdot I=I$.
@@
In brief, all exponential factors at infinity are shifted by $-\frac{\lambda}{2}z^{2}$ and all other exponential factors are left unchanged. As a consequence, the twist $T_{\lambda}$ preserves the global fission data and classes of global admissible deformations. The twist may change the slopes of the exponential factors at infinity, but leaves the slopes of Stokes circles at finite distance unchanged.
@@
Let $I$ be a Stokes circle such that $\pi(i)=a\in\mathbb{P}^{1}$, and $\breve{\mathcal{C}}_{I}\in\operatorname{GL}_{n}(\mathbb{C})$ a conjugacy class. Let $\breve{\mathcal{C}}_{S_{\lambda}\cdot I}\in\operatorname{GL}_{n}(\mathbb{C})$ be the conjugacy class such that $S_{\lambda}\cdot(I,\breve{\mathcal{C}}_{I})=(S_{\lambda}\cdot I,\breve{\mathcal{C}}_{T_{\lambda}\cdot I})$. For $\lambda\in\mathbb{C}^{*}$,$\pi(S_{\lambda}\cdot I)=a/\lambda$, and writing $I=\langle q(z_{a}^{-1})\rangle$ with $q\in\mathbb{C}[z^{-1/r}]$, we have $S_{\lambda}\cdot I=\langle q(\lambda z_{a/\lambda}^{-1})\rangle$. Furthermore $\breve{\mathcal{C}}_{S_{\lambda}\cdot I}=\breve{\mathcal{C}}_{I}$.
As a consequence, for any $\lambda\in\mathbb{C}$, the scaling $S_{\lambda}$ preserves the global fission data and classes of global admissible deformations, as well as the slopes of all Stokes circles.
@@
Let $(E,\nabla)$ be an irreducible connection on a Zariski open subset of $\mathbb{P}^{1}$ (different from a rank one connection with only a pole at infinity of order less than 2), and $A\in\operatorname{SL}_{2}(\mathbb{C})$. Then
$\Gamma(E,\nabla)=\Gamma(A\cdot(E,\nabla)).$
3. Nearby representations

### v3

@@ 2.3 Diagrams
    If $a_{I}=a_{J}=\infty$, then $B_{IJ}:=B^{\infty}_{IJ}$,
@@
    If $a_{I}=a_{J}\neq\infty$, then $B_{IJ}=B_{JI}:=B^{\infty}_{IJ}-\alpha_{I}\beta_{J}-\alpha_{J}\beta_{I}$,
@@
    If $a_{I}\neq\infty$ and $a_{J}=\infty$, then $B_{IJ}=B_{JI}:=\beta_{J}(\alpha_{I}+\beta_{I})$,
@@
    If $a_{I}\neq\infty,a_{J}\neq\infty$ and $a_{I}\neq a_{J}$, then $B_{IJ}=B_{JI}:=0$.
@@
The core diagram $\Gamma_{c}\bigl(\bm{\breve{\Theta}}\bigr)$ is entirely determined by the fission data of $\bm{\breve{\Theta}}$ together with the slopes of all Stokes circles at finite distance, because these data determine the numbers $A_{IJ}=\operatorname{irr}(\operatorname{Hom}(I,J))$}. In general, the diagram contains strictly less information, because $A_{IJ}$ is a sum of different contributions, coming from the levels of the common parts of IIJJ and the fission exponent $f_{I,J}$. As we shall review below, the interest of the (core) diagram is that, unlike the fission data, it is invariant under symplectic transformations.
2.4 Fourier transform and symplectic transformations
We now describe how the $\operatorname{SL}_{2}(\mathbb{C})$ transformations on irregular connections on $\mathbb{P}^{1}$ act on their formal data and fission data.
2.4.1 Fourier transform
[…]
@@ 2.4.1 Fourier transform
There exists a bijection, that we also denote by $F$, from the set of all modified formal data to itself, such that if $(E,\nabla)$ is a irreducible algebraic connection on a Zariski open subset of $\mathbb{P}^{1}$ which is not a rank one connection with a single singularity at infinity of Poincaré–Katz rank $\leq 1$, $\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$ its modified formal data, $(E^{\prime},\nabla^{\prime})$ is the Fourier transform of $(E,\nabla)$ and $\bigl(\bm{\breve{\Theta}^{\prime}},\bm{\breve{\mathcal{C}}}^{\prime}\bigr)$ its modified formal data, the following diagram commutes:
$(E,\nabla)$$(E^{\prime},\nabla^{\prime})$$\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$$\bigl(\bm{\breve{\Theta}^{\prime}},\bm{\breve{\mathcal{C}}}^{\prime}\bigr)$$\scriptstyle{F}$$\scriptstyle{F}$
This map $F$ is the formal Fourier transform. It acts independently on the Stokes circles: if we use the notation $\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)=\sum_{I\in N}\bigl(I,\breve{\mathcal{C}}_{I}\bigr)$ where $N$ is the set of active circles of $\bm{\breve{\Theta}}$, we have
$F\cdot\sum_{I\in N}\bigl(I,\breve{\mathcal{C}}_{I}\bigr)=\sum_{I\in N}F\cdot\bigl(I,\breve{\mathcal{C}}_{I}\bigr).$
@@
This bijection is actually given by a Legendre transform. It can be computed explicitly, and takes slightly different forms depending on whether the Stokes circles are at finite distance or at infinity, and in the latter case on whether its slope is greater or less than 1, see [20, Section 3.1]. The next proposition describes the ramification orders, slopes and levels of the formal Fourier transform $F\cdot I$ of any Stokes circle $I$ as a function of those of $I$, as well as the transformation of the conjugacy classes of formal monodromies.
@@
Let $I$ be a Stokes circle, and $\breve{\mathcal{C}}_{I}\in\operatorname{GL}_{n}(\mathbb{C})$ a conjugacy class. Let $\breve{\mathcal{C}}_{F\cdot I}\in\operatorname{GL}_{n}(\mathbb{C})$ be the conjugacy class such that $F\cdot\bigl(I,\breve{\mathcal{C}}_{I}\bigr)=\bigl(F\cdot I,\breve{\mathcal{C}}_{F\cdot I}\bigr)$. We have the following:
@@ 1. $(1)$
    If $I$ is of the form $\bigl\langle\alpha z\bigr\rangle_{\infty}$, with $\alpha\in\mathbb{C}$, then $F\cdot I=\bigl\langle 0\bigr\rangle_{\alpha}$, and $\breve{\mathcal{C}}_{F\cdot I}=\breve{\mathcal{C}}_{I}^{-1}$.
@@ 2. $(2)$
    If $I$ is of slope $\leq 1$ at infinity, of the form $\bigl\langle\alpha z+q\bigr\rangle_{\infty}$, with $\alpha\in\mathbb{C}$, and $q\neq 0$ of slope <1<1, then $F\cdot I$} is of the form $F\cdot I=\bigl\langle\widetilde{q}\bigr\rangle_{\alpha}$, with $\operatorname{irr}(\tilde{q})=s,\operatorname{ram}(\tilde{q})=r-s$, $\operatorname{Levels}(\tilde{q})=\frac{r}{r-s}\operatorname{Levels}(q)$. Also $\breve{\mathcal{C}}_{F\cdot I}=(-1)^{s}\breve{\mathcal{C}}_{I}^{-1}$.
@@ 3. $(3)$
    If $I$ is of slope >1>1 at infinity, with $\operatorname{ram}(I)=r$, $\operatorname{irr}(I)=s$, then $F\cdot I$ satisfies $\operatorname{ram}(F\cdot I)=s-r$, $\operatorname{irr}(F\cdot I)=s$, and $\operatorname{Levels}(F\cdot I)=\frac{r}{s-r}\operatorname{Levels}(I)$. Also $\breve{\mathcal{C}}_{F\cdot I}=(-1)^{s}\breve{\mathcal{C}}_{I}$.
@@ 4. $(4)$
    If $I=\bigl\langle q\bigr\rangle_{a}$ is an irregular circle at finite distance for $a\in\mathbb{C}=\mathbb{P}^{1}\smallsetminus\{\infty\}$, with $q\neq 0$}, $\operatorname{ram}(q)=r$, and $\operatorname{irr}(q)=s$, then $F\cdot I$ is a circle of slope $\leq 1$, at infinity, of the form $\bigl\langle-az+\tilde{q}\bigr\rangle_{\infty}$}, with $\operatorname{ram}(\tilde{q})=r+s,\operatorname{irr}(\tilde{q})=s$, and $\operatorname{Levels}(\tilde{q})=\frac{r}{r+s}\operatorname{Levels}(q)$. Also $\breve{\mathcal{C}}_{F\cdot I}=(-1)^{s}\breve{\mathcal{C}}_{I}^{-1}$.
@@ 5. $(5)$
    If $I=\bigl\langle 0\bigr\rangle_{a}$ is a tame circle at finite distance, then $F\cdot I=\bigl\langle-az\bigr\rangle_{\infty}$, and $\breve{\mathcal{C}}_{F\cdot I}=\breve{\mathcal{C}}_{I}^{-1}$.
@@
All statements except those describing the sets of levels are already contained in the works [41, 47] establishing the stationary phase formula. For the levels, some of the statements are proven in [32] using results on singularities of curves. Let us discuss how to derive them all more directly from results in [20]. In the case (3) where $I$ is at infinity and of slope >1>1, the statement for the levels can be deduced quite directly from [20, Proposition 4.8 and the proof of Theorem 4.9], as follows. Proposition 4.8 says that all exponents of $F\cdot I$ belong to a certain set $K\subset\mathbb{Q}_{>0}$} determined explicitly from $I$. Next, although it does not use the terminology of levels, the proof of Theorem 4.9 shows that, inside $K$, the subset $\operatorname{Levels}(F\cdot I)$ is exactly $\frac{r}{s-r}\operatorname{Levels}(I)$}. This is because, in the explicit formula for the number of edge-loops $B_{I,I}$ of a Stokes circle $I$, given by Lemma 4.6 of loc. cit., the nonzero positive terms are in one-to-one correspondence with the levels of $I$ (by definition of the notion of levels). The proof of Theorem 4.9 establishes that there is a one-to-one correspondence between the nonzero positive terms in $B_{I,I}$ and $B_{F\cdot I,F\cdot I}$, i.e., there is a bijection between $\operatorname{Levels}(I)$, and $\operatorname{Levels}(F\cdot I)$, given by $\operatorname{Levels}(F\cdot I)=\frac{r}{s-r}\operatorname{Levels}(I)$.
The other cases are obtained in a completely similar way, by using the expressions for the Legendre transform for the different cases given, e.g., in [20, Section 3.1]. The point is that the difference between the cases concerns only the transformation of the leading terms of the Stokes circles, i.e., the change of slopes, without affecting what happens for the subleading terms. ∎
To fully determine the fission data of the Fourier transform, we must also describe the fission exponents between distinct Stokes circles of the Fourier transform which are at the same singularities. This can be done explicitly:
@@
Let $I$, $J$ be two distinct Stokes circles. There are the following possibilities:
@@
    • If $I$ and $J$ have a common part, denoting by $k$ its slope, we have $f_{F\cdot I,F\cdot J}=\frac{k}{k+1}f_{I,J}$.
@@
    If $\pi(I)=\pi(J)=\infty$, $I,J$ are both of slope $\leq 1$ and are of the form $I=\bigl\langle\alpha z+q\bigr\rangle$, $J=\bigl\langle\alpha z+q^{\prime}\bigr\rangle$ with $\alpha\in\mathbb{C}$, and $q$, $q^{\prime}$ of slope <1<1, then $\pi(F\cdot I)=\pi(F\cdot J)=\alpha$ and
    • If $q$ and $q^{\prime}$ have no common part and respective slopes $k$, $l$, then $f_{F\cdot I,F\cdot J}=\max\bigl(\frac{k}{1-k},\allowbreak\frac{l}{1-l}\bigr)$.
    • If $q$ and $q^{\prime}$ have a common part $q_{c}$, denoting by $k$ its slope, we have $f_{F\cdot I,F\cdot J}=\frac{k}{1-k}f_{q,q^{\prime}}$.
@@
    If $\pi(I)=\pi(J)=\infty$, and $I$, $J$ are both of slope >1>1, then $\pi(F\cdot I)=\pi(F\cdot J)=\infty$ and
    • If $I$ and $J$ have no common part, then denoting $k=\operatorname{slope}(I)$, $l=\operatorname{slope}(J)$, we have $f_{I,J}=\max\bigl(\frac{k}{k-1},\frac{l}{l-1}\bigr)$.
    • If $I$ and $J$ have a common part, let $k$ denote its slope, then $f_{F\cdot I,F\cdot J}=\frac{k}{k-1}f_{I,J}$.
@@
All statements where there is no common part involved follow directly from the previous proposition. For the statements in the cases where there is a common part, in the case where IIJJ are at infinity and of slope >1>1 this is established in the proof of [20, Theorem 5.9] (the word ‘fission exponent’ was not used there: in terms of the terminology used there it corresponds to the highest number among the exponents of the leading term of the different parts of $F\cdot I$ and $F\cdot J$). The other cases are similar, using again the expressions for the Legendre transform for the different cases given in [20, Section 3.1], and noticing that this just changes the transformation of the slopes, without affecting the subleading terms (hence in particular the fission exponents when the common part is nonzero). ∎
2.4.2 General symplectic $\bm{\operatorname{SL}_{2}(\mathbb{C})}$ transformations
The Fourier–Laplace transform is part of a larger group of transformations acting on modules over the Weyl algebra. Indeed, to any matrix $A=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)$ in $\operatorname{SL}_{2}(\mathbb{C})$, we can associate an automorphism of the Weyl algebra $A_{1}=\mathbb{C}[z]\bigl\langle\partial_{z}\bigr\rangle$ given by $z\mapsto az+b\partial_{z},\partial_{z}\mapsto cz+d\partial_{z}$. This induces an action of the group $\operatorname{SL}_{2}(\mathbb{C})$ of symplectic transformations on modules over the Weyl algebra (see [41, p. 87]).
@@ 2.4.2 General symplectic $\bm{\operatorname{SL}_{2}(\mathbb{C})}$ transformations
    the Fourier–Laplace transform $F$, corresponding to the matrix $\left(\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right)$;
@@
    twists at infinity $T_{\lambda}$, for $\lambda\in\mathbb{C}$, corresponding to the matrix $\left(\begin{smallmatrix}1&\lambda\\ 0&1\end{smallmatrix}\right)$;
@@
    scalings $S_{\lambda}$, for $\lambda\in\mathbb{C}^{*}$, corresponding to the matrix $\left(\begin{smallmatrix}\lambda^{-1}&0\\ 0&\lambda\end{smallmatrix}\right)$.
The geometric interpretation of twists and scalings is the following. The twist $T_{\lambda}$ corresponds to taking the tensor product with the rank one module $(\mathbb{C}[z],\partial_{z}+\lambda z)$, and the scaling $S_{\lambda}$ corresponds to perform the change of variable $z\mapsto z/\lambda$ on $\mathbb{P}^{1}$.
As for the Fourier transform, any element of $\operatorname{SL}_{2}(\mathbb{C})$ induces a transformation on modules over the Weyl algebra, and on irreducible connections on Zariski open subsets of the affine line. Moreover, there exists again a corresponding formal transformation (that we will also denote by $A$) on modified formal data with a similar structure. The action of twists and scalings on formal data and fission data is much simpler than for the Fourier transform.
@@
Let $I$ be a Stokes circle, and $\breve{\mathcal{C}}_{I}\in\operatorname{GL}_{n}(\mathbb{C})$ a conjugacy class. Let $\breve{\mathcal{C}}_{T_{\lambda}\cdot I}\in\operatorname{GL}_{n}(\mathbb{C})$ be the conjugacy class such that $T_{\lambda}\cdot\bigl(I,\breve{\mathcal{C}}_{I}\bigr)=(T_{\lambda}\cdot I,\breve{\mathcal{C}}_{T_{\lambda}\cdot I})$. We have the following:
@@
    If $\pi(I)=\infty$, then $\pi(T_{\lambda}\cdot I)=\infty$, and if $I=\bigl\langle q\bigr\rangle_{\infty}$, we have $T_{\lambda}\cdot I=\bigl\langle q-\frac{\lambda}{2}z^{2}\bigr\rangle_{\infty}$. In particular, $T_{\lambda}\cdot I$ has the same ramification order and set of levels as $I$, $\operatorname{slope}(T_{\lambda}\cdot I)=\max(\operatorname{slope}(\bigl\langle q\bigr\rangle),\operatorname{slope}\bigl(\bigl\langle q-\frac{\lambda}{2}z^{2}\bigr\rangle\bigr)$, for $\lambda\in\mathbb{C}$.
@@
    Otherwise, if $\pi(I)\neq\infty$, $T_{\lambda}\cdot I=I$.
@@
In brief, all exponential factors at infinity are shifted by $\frac{\lambda}{2}z^{2}$ and all other exponential factors are left unchanged. As a consequence, the twist $T_{\lambda}$ preserves the global fission data and classes of global admissible deformations. The twist may change the slopes of the exponential factors at infinity, but leaves the slopes of Stokes circles at finite distance unchanged.
@@
Let $I$ be a Stokes circle such that $\pi(i)=a\in\mathbb{P}^{1}$, and $\breve{\mathcal{C}}_{I}\in\operatorname{GL}_{n}(\mathbb{C})$ a conjugacy class. Let $\breve{\mathcal{C}}_{S_{\lambda}\cdot I}\in\operatorname{GL}_{n}(\mathbb{C})$ be the conjugacy class such that $S_{\lambda}\cdot\bigl(I,\breve{\mathcal{C}}_{I}\bigr)=\bigl(S_{\lambda}\cdot I,\breve{\mathcal{C}}_{T_{\lambda}\cdot I}\bigr)$. For $\lambda\in\mathbb{C}^{*}$, $\pi(S_{\lambda}\cdot I)=a/\lambda$, and writing $I=\bigl\langle q\bigl(z_{a}^{-1}\bigr)\bigr\rangle$ with $q\in\mathbb{C}\big[z_{a}^{-1/r}\big]$, using the local coordinate $z_{a}:=z-a$ if $a\neq\infty$, and $z_{\infty}:=z^{-1}$ if $a=\infty$, we have $S_{\lambda}\cdot I=\bigl\langle q\bigl(\lambda z_{a/\lambda}^{-1}\bigr)\bigr\rangle$. Furthermore $\breve{\mathcal{C}}_{S_{\lambda}\cdot I}=\breve{\mathcal{C}}_{I}$.
As a consequence, for any $\lambda\in\mathbb{C}^{*}$, the scaling $S_{\lambda}$ preserves the global fission data and classes of global admissible deformations, as well as the slopes of all Stokes circles.
@@
Let $(E,\nabla)$ be an irreducible connection on a Zariski open subset of $\mathbb{P}^{1}$ (different from a rank one connection with only a pole at infinity of order less than $2)$, and $A\in\operatorname{SL}_{2}(\mathbb{C})$. Then $\Gamma(E,\nabla)=\Gamma(A\cdot(E,\nabla))$.
3 Classes of nearby representations